A satellite was launched in 1983. It followed an elliptical orbit that varied from about 2.00 x 10^3 km to 5.00 x 10^3 km above earth. What was its orbital speed at its closest approach to earth?


1
Expert's answer
2021-12-17T11:50:02-0500

Explanations & Calculations


  • If the satellite was launched at the highest margin: "\small 5000\,km" in 1983, then it was meant to stay in that orbit (ideally) moving at the corresponding speed of that orbit.
  • That was,

"\qquad\qquad
\begin{aligned}
\small G\frac{M_Em}{r_1^2}&=\small m\frac{v_1^2}{r_1}\\
\small v_1&=\small \sqrt{\frac{GM_E}{r_1}}
\end{aligned}"


  • When it approaches the lowest orbit, its speed is meant to increase according to the law of conservation of angular momentum ("\small I=mvr" ).
  • Then, it can be calculated as follows.

"\qquad\qquad
\begin{aligned}
\small mv_1r_1&=\small mv_2r_2\\
\small v_2&=\small v_1.\frac{r_1}{r_2}\\
&=\small \sqrt{\frac{GM_E}{r_1}}.\Big(\frac{r_1}{r_2}\Big)
\end{aligned}"

  • Here G and "\small M_E" are the Universal gravitational constant and the mass of the earth respectively.